OpenStudy (anonymous):

How many solutions are there to the following system of equations? 5x + 3y = 7 2x – y = 2 A. 2 B. 0 C. infinitely many D. 1

OpenStudy (anonymous):

please help

hero (hero):

Hint 2x - 2 = y

hero (hero):

If there are two lines, there will only either be 0, 1, or infinite solutions. So you can eliminate A as a possible choice.

OpenStudy (***[isuru]***):

Hi, 1st Number the 2 equation 5x + 3y = 7 ---(1) 2x – y = 2 ---(2) multiply (2) by 3 and add it to (1) u could eliminate y and find x then use substitute

hero (hero):

If the equations are not the same, then we can eliminate infinite solutions. If both equations have different slopes, then we can eliminate zero solutions.

hero (hero):

Attempting to solve the system, however, reveals exactly how many solutions there are.

hero (hero):

So using the expression for y, which is 2x - 2, substitute that into the first equation to get: 5x + 3(2x - 2) = 7 And solve for x. If you are able to, then you'll know how many solutions exist.

OpenStudy (anonymous):

oh ok thnx